# Distillation versus Contrastive Learning: How to Train Your Rerankers

### v1

@@ Abstract
Training text rerankers is crucial for information retrieval. Two primary strategies are widely used: contrastive learning (optimizing directly on ground-truth labels) and knowledge distillation (transferring knowledge from a larger reranker). While both have been studied in the literature, a clear comparison of their effectiveness for training cross-encoder rerankers under practical conditions is needed.
This paper empirically compares these strategies by training rerankers of different sizes and architectures using both methods on the same data, with a strong contrastive learning model acting as the distillation teacher. Our results show that knowledge distillation generally yields better in-domain and out-of-domain ranking performance than contrastive learning when distilling from a larger teacher model. This finding is consistent across student model sizes and architectures. However, distilling from a teacher of the same capacity does not provide the same advantage, particularly for out-of-domain tasks. These findings offer practical guidance for choosing a training strategy based on available teacher models. Therefore, we recommend using knowledge distillation to train smaller rerankers if a larger, more powerful teacher is accessible; in its absence, contrastive learning provides a strong and more reliable alternative otherwise.
@@
Zhichao Xu 12 Zhiqi Huang 3 Shengyao Zhuang 4 Ashim Gupta 1 Vivek Srikumar 1 1 Kahlert School of Computing, University of Utah 2 Scientific Computing and Imaging Institute, University of Utah 3 Capital One, Inc    4 The University of Queensland zhichao.xu@utah.edu    svivek@cs.utah.edu
@@ 1 Introduction
Modern information retrieval (IR) systems often rely on a two-stage process: an initial retriever quickly finds candidate texts, and a more powerful reranker re-orders them by relevance to improve results (Schütze et al., 2008; Ma et al., 2023; Asai et al., 2024; Singh et al., 2025, inter alia). Reranking involves scoring and sorting texts based on a given query, and cross-encoder models are commonly used for this task Yates et al. (2021); Zhuang et al. (2024b).
@@
Effective training is key to building a high-performing cross-encoder reranker. Two main strategies stand out in the field. The first is contrastive learning (CL), where the model is trained directly using ground-truth relevance labels, learning to distinguish positive (relevant) examples from negative (irrelevant) ones Oord et al. (2018); Gao et al. (2021); Ma et al. (2023). The second is knowledge distillation (KD), which involves training a smaller “student” model to replicate the outputs of a larger, more capable “teacher” model Hinton et al. (2015); Gou et al. (2021); Hofstätter et al. (2020); Schlatt et al. (2025). This method is often employed to create efficient models that can approximate the performance of larger ones.
@@
While both strategies are used in practice to train rerankers, a direct, controlled empirical comparison of their effectiveness specifically for training cross-encoder rerankers is missing from literature. Hofstätter et al. (2020) explored distilling from rerankers to retrievers; while Baldelli et al. (2024); Schlatt et al. (2025) focused on training rerankers via distilling from large language models (LLMs) like GPT-4 Achiam et al. (2023).
@@
Seeing this gap, the objective of this paper is to empirically compare contrastive learning and knowledge distillation as training strategies for cross-encoder rerankers. To achieve this objective, we conduct a controlled experiment. We trained cross-encoder models of various sizes (0.5B, 1.5B, 3B, 7B) and different architectures (Transformer and Recurrent) using both contrastive learning and knowledge distillation. For knowledge distillation, we employ a powerful 7B Qwen2.5 model Yang et al. (2024) trained with contrastive learning as the teacher. We did not consider smaller encoder models like BERT Devlin et al. (2019), as recent works have reported decoder-only models’ effectiveness (Ma et al., 2023; Muennighoff et al., 2024a; Zhang et al., 2025, inter alia). We train all models on the same dataset and evaluated their ranking performance on standard in-domain and out-of-domain benchmarks.
Our experimental results consistently show that training with knowledge distillation yields better ranking performance than direct contrastive learning when the student model is smaller than the teacher model. This performance gain was observed on both in-domain and out-of-domain evaluation datasets and across different student model sizes and architectures. For instance, distilling from our 7B teacher improved the performance of 0.5B, 1.5B, and 3B student models compared to training them with contrastive learning alone. However, we also found that distilling from a teacher of the same capacity (7B student distilling from 7B teacher) did not necessarily lead to performance improvements, particularly impacting out-of-domain generalization, hinting the training strategy’s limitation Gou et al. (2021); Gholami and Omar (2023). Together, these findings suggest that knowledge distillation is an effective strategy for training smaller rerankers when a strong, larger teacher model is available, but its benefits may diminish when the student capacity approaches or matches the teacher’s, especially concerning out-of-domain robustness.
@@ 2 Methodology
In the rest of this section, we first formally define the text reranking problem and our notation (§ 2.1). We then provide detailed technical descriptions of the training process using contrastive learning (§ 2.2) and knowledge distillation (§ 2.3).
@@ 2.1 The Text Reranking Problem
Modern IR systems often employ a two-stage retrieval-and-rerank pipeline (Schütze et al., 2008; Zhang et al., 2021b; Asai et al., 2024, inter alia). An efficient first-stage retriever initially fetches a broad set of candidate texts. Subsequently, a more powerful reranker refines this initial list to optimize ranking metrics. Reranking is the task of ordering texts (e.g., passages or documents) by their relevance to a given query.
@@ 2.2 Training with Contrastive Learning
Contrastive learning is the prevalent training strategy for learning representations Gutmann and Hyvärinen (2010); Oord et al. (2018); Weng (2021). To understand its application, we first review the underlying principle of the InfoNCE loss Oord et al. (2018), which is derived from Noise-Contrastive Estimation.
The general goal is to learn a model that distinguishes a “positive” data sample from a set of “negative” (or noise) samples, given a certain context. Denote a context vector $\mathbf{c}$, and consider a set of $N$ samples $X=\{\mathbf{x}_{i}\}_{i=1}^{N}$, where one sample $\mathbf{x}_{\text{pos}}$ is a positive sample drawn from the condictional distribution $p(\mathbf{x}|\mathbf{c})$, and the $N-1$ negative samples are drawn from a proposal distribution $p(\mathbf{x})$. Using Bayes’ rule, the probability that a sample $\mathbf{x}_{i}$ is the positive one is:
@@
We can define the scoring function that is propotional to the density ratio $f(\mathbf{x},\mathbf{c})\propto\frac{p(\mathbf{x},\mathbf{c})}{p(\mathbf{c})}$. The InfoNCE loss optimizes the negative log probability of classifying the positive sample correctly:
@@
Applying to text reranking, we map the abstract concepts to our task:
@@ 2.3 Training with Knowledge Distillation
$\mathcal{L}_{\text{KD}}=D_{\mathrm{KL}}\left(\mathrm{softmax}\left(\frac{\mathbf{z}_{t}}{T}\right)\bigg{|}\mathrm{softmax}\left(\frac{\mathbf{z}_{s}}{T}\right)\right)$ (3)
where $T$ is the temperature hyperparameter. A higher temperature creates a softer probability distribution, which can help in transferring more nuanced information from the teacher. In practice, $T$ is often set to 1 Hinton et al. (2015); Zhang et al. (2021a).
Contrastive learning directly optimizes a model on ground-truth relevance labels, whereas knowledge distillation leverages the nuanced, soft-labeled outputs of a more capable teacher model. Having detailed these two distinct training paradigms, the central goal of this paper is to conduct a direct empirical comparison of their effectiveness for training cross-encoder rerankers, under a controlled experimental setting. We will detail our experimental setup in § 3.
@@ 3 Experimental Setup
To construct the training set for the rerankers, we follow the same strategy as Ma et al. (2023) to mine hard negatives from the first stage retriever. Specifically, for each $(q_{i},d_{i}^{+})$ pair, we randomly sample $k$ negatives from top-200 passages retrieved by the retriever, excluding $d_{i}^{+}$. As shown in the experimental results (§ 4), the models trained using this constructed dataset achieves on par or improved performance over RankLlama2 Ma et al. (2023), suggesting the correctness of the pipeline. We did not further investigate hard negative mining strategies like prior works Yu et al. (2024); Thakur et al. (2025); Lee et al. (2025) as they are orthogonal to the goal of this paper.
@@
We evaluate the reranking performance following prior works’ practices (Ma et al., 2023; Xu et al., 2025c, inter alia). For in-domain evaluation, we use MS MARCO passage dev set, which includes 6,980 queries. We also include TREC DL19 and DL20 Craswell et al. (2020, 2021) consisting of 43 and 54 queries respectively. For out-of-domain evaluation, we adopt BEIR benchmark Thakur et al. (2021). We evaluate performances on 13 subsets that are publicly available. Refer to Appx. A for details of datasets.
@@
Our main experiments are based on RepQwen — retriever with Qwen2.5-7B backbone following RepLlama Ma et al. (2023) training strategy. We also train a retriever with Llama-3.1-8B backbone and report its performance.
For our main focus of study — reranker models — we use Qwen2.5 models of different sizes, including 0.5B, 1.3B, 3B and 7B, to allow us to observe the scaling trend. For each model size, we train two reranker models: one via direct contrastive learning and one via distillation; we denote with CL/KD suffix, e.g., RankQwen-0.5B-CL means 0.5B model trained with contrastive learning. For knowledge distillation training, we distil from the RankQwen-7B-CL teacher.
We also experiment with RecurrentGemma De et al. (2024), a recurrent language model based on the Griffin architecture instead of quadratic complexity Transformers Vaswani et al. (2017). We use the 2B variant and refer to the trained model as RankRGemma-2B. Recent works have explored recurrent language models’ efficacy for IR tasks, such as state space models like Mamba Gu and Dao (2023); Dao and Gu (2024) for retrieval Zhang et al. (2024) and reranking Xu (2024); Xu et al. (2025c). We followup along this direction by examine recurrent language models’ efficacy under different training strategies.
@@
We include a few prior works’ results under the same training setting. RepLlama and RankLlama Ma et al. (2023) are the closest baselines which use the same trainset for contrastive learning, using Llama-2 backbone Touvron et al. (2023). CSPLADE Xu et al. (2025a) is a learned sparse retrieval model with Llama-3-8B backbone Dubey et al. (2024), and achieves competitive performance compared to dense retrieval models. RankMamba-2 Xu et al. (2025c) trains Mamba-2-based rerankers for passage reranking, though their results are based on BGE retriever Xiao et al. (2023). We use the original numbers reported by the authors.
@@
Our implementation is based on packages including PyTorch, Huggingface Transformers and Tevatron-v2 Ma et al. (2025). For all our models, we train with LoRA Hu et al. (2021) to balance in-domain and out-of-domain performance and reduce overfitting. We apply techniques including DeepSpeed stage 2 Aminabadi et al. (2022), activation checkpointing, FlashAttention-2 Dao (2024), mixed precision and gradient accumulation for scalable training.
Hyperparameters.
For all our training runs, we use the same set of optimized hyperparameters identified from prior works and our preliminary experiments Ma et al. (2023); Xu et al. (2025a). We use LoRA rank=16, peak learning rate=1e-4 with linear warmup. We use 8 GPUs with per device batch 2, gradient accumulation steps 4, which leads to a global batch size of 64. For each $(q_{i},d_{i}^{+})$ pair, we use $k=15$ negatives as we find increasing to $31$ negatives may lead to instable contrastive learning. For RepQwen training, we use in-batch negatives; while for reranker training, we focus solely on each $(q_{i},d_{i}^{+})$ pair’s own hard negatives.
@@ 4 Results and Analysis
cross-SimLM Wang et al. (2022) 110M bi-SimLM 200 43.7 74.6 72.7
@@
In Table 1, we report the reranking performance on MS MARCO Dev and DL19+DL20. We observe our trained retrieval models, i.e., RepLlama3 and RepQwen improves on par or better performance compared to RepLlama2 and CSPLADE, two models trained with same trainset. Similarly, RankQwen-7B-CL is comparable to RankLlama2-7B, suggesting the correctness of our training pipeline.
@@
We also note RankQwen-7B-KD achieves similar in-domain performance as RankQwen-3B-KD and the RankQwen-7B-CL teacher, which suggests that the student model is not benefiting from knowledge distillation training when the teacher model is of the same capacity. We will revisit this problem in out-of-domain results.
@@
We report the out-of-domain evaluation results in Table 2. The baseline methods’ results are deferred to Appx. B.
We note the similar observations as the in-domain evaluation. Directly comparing contrastive learning versus knowledge distillation, RankQwen-0.5B-KD achieves similar performance as RankQwen-0.5B-CL (53.1 average NDCG@10 v.s. 53.3); while RankQwen-1.5B-KD, RankQwen-3B-KD, RankRGemma-2B have better performance over their contrastive learning counterparts.
@@
Scaling trend.
Scaling has been proven effective in IR tasks (Neelakantan et al., 2022; Muennighoff et al., 2024b; Xu et al., 2025b; Zhu et al., 2023, inter alia). We also compare the scaling trend between CL versus KD for passage reranking in Fig. 1. We make two observations:

### v2

@@ Abstract
Training effective text rerankers is crucial for information retrieval. Two strategies are widely used: contrastive learning (optimizing directly on ground-truth labels) and knowledge distillation (transferring knowledge from a larger reranker). While both have been studied extensively, a clear comparison of their effectiveness for training cross-encoder rerankers under practical conditions is needed.
This paper empirically compares these strategies by training rerankers of different sizes and architectures using both methods on the same data, with a strong contrastive learning model acting as the distillation teacher. Our results show that knowledge distillation generally yields better in-domain and out-of-domain ranking performance than contrastive learning when distilling from a larger teacher model. This finding is consistent across student model sizes and architectures. However, distilling from a teacher of the same capacity does not provide the same advantage, particularly for out-of-domain tasks. These findings offer practical guidance for choosing a training strategy based on available teacher models. We recommend using knowledge distillation to train smaller rerankers if a larger, more powerful teacher is accessible; in its absence, contrastive learning remains a robust baseline.
@@
Zhichao Xu 12 Zhiqi Huang 3 Shengyao Zhuang 4 Vivek Srikumar 1 1 Kahlert School of Computing, University of Utah 2 Scientific Computing and Imaging Institute, University of Utah 3 Capital One, Inc    4 The University of Queensland zhichao.xu@utah.edu    svivek@cs.utah.edu
@@ 1 Introduction
Modern information retrieval (IR) systems often rely on a two-stage process: an initial retriever quickly finds candidate texts, and a more powerful reranker re-orders them by relevance to improve the final ranking quality (Schütze et al., 2008; Ma et al., 2023; Asai et al., 2024; Singh et al., 2025, inter alia). This reranking stage, which involves scoring and sorting texts by their relevance to a given query, commonly employs powerful cross-encoders Yates et al. (2021); Zhuang et al. (2024b).
@@
Effective training is key to building a high-performing cross-encoder reranker. Two main strategies have emerged for this purpose. The first, contrastive learning (CL), trains the model directly using ground-truth relevance labels, learning to distinguish positive (relevant) examples from negative (irrelevant) ones Oord et al. (2018); Gao et al. (2021); Ma et al. (2023). The second, knowledge distillation (KD), instead involves training a smaller “student” model to replicate the outputs of a larger, more capable “teacher” model Buciluǎ et al. (2006); Hinton et al. (2015); Hofstätter et al. (2020); Schlatt et al. (2025). This method is often employed to create efficient models that can approximate the performance of larger ones.
@@
Both strategies are widely used to train rerankers. But for practical deployments, is one preferable to the other? Prior work has explored related, but distinct, questions. For instance, Hofstätter et al. (2020) explored distilling from rerankers to retrievers; while Baldelli et al. (2024) and Schlatt et al. (2025) focused on training rerankers via distilling from large language models (LLMs) like GPT-4 Achiam et al. (2023). However, there is no direct comparison of these two strategies.
@@
In this paper, we address this gap. We present an empirical comparison of contrastive learning and knowledge distillation as training strategies for cross-encoder rerankers. To this end, we conduct a suite of controlled experiments. We train cross-encoder models of various sizes (0.5B, 1.5B, 3B, 7B) and different architectures (Transformer and Recurrent) using both strategies. For knowledge distillation, we employ a performant 7B Qwen2.5 model Yang et al. (2024) trained with contrastive learning as the teacher. We focus on modern decoder-only models, as recent works have reported their effectiveness over smaller encoder-based models like BERT (Ma et al., 2023; Muennighoff et al., 2024a; Zhang et al., 2025, inter alia). We train all models on the same dataset and evaluate their ranking performance on standard in-domain and out-of-domain benchmarks.
Our findings consistently show that training with knowledge distillation yields better ranking performance than direct contrastive learning when the student model is smaller than the teacher. For instance, distilling from our 7B teacher significantly improved the performance of 0.5B, 1.5B, and 3B student models on both in-domain and out-of-domain benchmarks compared to training them with contrastive learning. This advantage holds across both Transformer and Recurrent model architectures. However, this benefit diminishes when the student and teacher have the same capacity; distilling from a 7B teacher to a 7B student yielded no significant improvements and, in some cases, harmed out-of-domain generalization Gou et al. (2021); Gholami and Omar (2023). To further test the robustness of our conclusions, we conduct an additional experiment where models are trained on a diverse, multi-domain dataset RLHN (Thakur et al., 2025) instead of only on hard negatives mined from a single source. In this setting, knowledge distillation again proves to be the more effective strategy, confirming its advantage even when the reranker’s training data is decoupled from the first-stage retriever. Together, these findings position knowledge distillation as a powerful and robust strategy for training smaller rerankers, provided a strong, larger teacher is available. In its absence, we find contrastive learning remains a strong baseline.
@@ 2 Methodology
In the rest of this section, we first formally define the text reranking problem and our notation (§˜2.1). We then provide detailed technical descriptions of the training process using contrastive learning (§˜2.2) and knowledge distillation (§˜2.3).
@@ 2.1 The Text Reranking Problem
Modern IR systems often employ a two-stage retrieval-and-rerank pipeline (Schütze et al., 2008; Zhang et al., 2021; Asai et al., 2024, inter alia). An efficient first-stage retriever initially fetches a broad set of candidate texts. Subsequently, a more powerful reranker refines this initial list to optimize ranking metrics. Reranking is the task of ordering texts (e.g., passages or documents) by their relevance to a given query.
@@ 2.2 Training with Contrastive Learning
Contrastive learning is one strategy for learning representations Oord et al. (2018); Weng (2021). Its application to reranking builds on the principle of the InfoNCE loss Oord et al. (2018), which is derived from Noise-Contrastive Estimation Gutmann and Hyvärinen (2010).
The general goal is to learn a model that distinguishes a “positive” data sample from a set of “negative” (or noise) samples, given a certain context. Denote a context vector $\mathbf{c}$, and consider a set of $N$ samples $X=\{\mathbf{x}_{i}\}_{i=1}^{N}$, where one sample $\mathbf{x}_{\text{pos}}$ is a positive sample drawn from the conditional distribution $p(\mathbf{x}|\mathbf{c})$, and the $N-1$ negative samples are drawn from a proposal distribution $p(\mathbf{x})$. Using Bayes’ rule, the probability that a sample $\mathbf{x}_{i}$ is the positive one is:
@@
We can define the scoring function that is proportional to the density ratio $f(\mathbf{x},\mathbf{c})\propto\frac{p(\mathbf{x},\mathbf{c})}{p(\mathbf{c})}$. The InfoNCE loss optimizes the negative log probability of classifying the positive sample correctly:
@@
We map the abstract concepts to our reranking task:
@@ 2.3 Training with Knowledge Distillation
$\mathcal{L}_{\text{KD}}=D_{\mathrm{KL}}\left(\mathrm{softmax}\left(\frac{\mathbf{z}_{t}}{T}\right)\bigg{|}\mathrm{softmax}\left(\frac{\mathbf{z}_{s}}{T}\right)\right)$
where $T$ is the temperature hyperparameter. A higher temperature creates a softer probability distribution, which can help in transferring more nuanced information from the teacher. In practice, $T$ is often set to 1 Hinton et al. (2015).
Whereas contrastive learning optimizes a model on ground-truth labels, knowledge distillation mimics the outputs of a more capable teacher. The central goal of this paper is to empirically compare these distinct paradigms for training cross-encoder rerankers under controlled settings, which we detail in the following section.
@@ 3 Experimental Setup
To construct the training set for the rerankers, we follow the same strategy as Ma et al. (2023) to mine hard negatives from the first stage retriever. Specifically, for each $(q_{i},d_{i}^{+})$ pair, we randomly sample $k$ negatives from top-200 passages retrieved by the retriever, excluding $d_{i}^{+}$. As shown in the experimental results (§˜4), the models trained using this constructed dataset achieves on par or improved performance over RankLlama2 Ma et al. (2023), suggesting the correctness of the pipeline. We did not further investigate hard negative mining strategies like prior works Yu et al. (2024); Thakur et al. (2025); Lee et al. (2025) as they are orthogonal to the goal of this paper.
@@
We evaluate the reranking performance following prior works’ practices (Ma et al., 2023; Xu et al., 2025c, inter alia). For in-domain evaluation, we use MS MARCO passage dev set, which includes 6,980 queries. We also include TREC DL19 and DL20 Craswell et al. (2020, 2021) consisting of 43 and 54 queries respectively. For out-of-domain evaluation, we adopt BEIR benchmark Thakur et al. (2021). We evaluate performances on 13 subsets that are publicly available. Refer to Appx.˜A for details of datasets.
@@
Our main experiments are based on RepQwen, the retriever with a Qwen2.5-7B backbone following the RepLlama Ma et al. (2023) training strategy. We also report the performance of a retriever with a Llama-3.1-8B backbone.
The objects of our study are reranker models. We evaluate Qwen2.5 models of different sizes, including 0.5B, 1.3B, 3B and 7B, to allow us to observe the scaling trend. For each model size, we train two reranker models: one via direct contrastive learning and one via distillation; we denote with CL/KD suffix, e.g., RankQwen-0.5B-CL means 0.5B model trained with contrastive learning. For knowledge distillation training, we distil from the RankQwen-7B-CL teacher.
We also experiment with RecurrentGemma De et al. (2024), a recurrent language model based on the Griffin architecture instead of quadratic complexity Transformers Vaswani et al. (2017). We use the 2B variant and refer to the trained model as RankRGemma-2B. Recent works have explored recurrent language models’ efficacy for IR tasks, such as state space models like Mamba Gu and Dao (2023); Dao and Gu (2024) for retrieval Zhang et al. (2024) and reranking Xu (2024); Xu et al. (2025c). We follow this direction to examine recurrent language models’ efficacy under different training strategies.
@@
Our baselines are prior results under the same training setting. RepLlama and RankLlama Ma et al. (2023) are the closest baselines which use the same trainset for contrastive learning, using Llama-2 backbone Touvron et al. (2023). CSPLADE Xu et al. (2025a) is a learned sparse retrieval model with Llama-3-8B backbone Dubey et al. (2024), and achieves competitive performance compared to dense retrieval models. RankMamba-2 Xu et al. (2025c) trains Mamba-2-based rerankers for passage reranking, though their results are based on BGE retriever Xiao et al. (2023). We compare against the original numbers reported by the authors.
@@
Our implementation is based on packages including PyTorch, Huggingface Transformers and Tevatron-v2 Ma et al. (2025). For all our models, we train with LoRA Hu et al. (2021) to balance in-domain and out-of-domain performance and reduce overfitting. For scalable training, we use DeepSpeed stage 2 Aminabadi et al. (2022), activation checkpointing, FlashAttention-2 Dao (2024), mixed precision and gradient accumulation. Appx.˜B gives details about hyperparameters.
@@ 4.1 Main Results
cross-SimLM Wang et al. (2022a) 110M bi-SimLM 200 43.7 74.6 72.7
@@
Table˜1 reports the reranking performance on MS MARCO Dev and DL19+DL20. We observe our trained retrieval models, i.e., RepLlama3 and RepQwen performs on par or better than RepLlama2 and CSPLADE, the two models trained with same trainset. Similarly, RankQwen-7B-CL is comparable to RankLlama2-7B, suggesting the correctness of our training pipeline.
@@
We also note RankQwen-7B-KD achieves similar in-domain performance as RankQwen-3B-KD and the RankQwen-7B-CL teacher, which suggests that the student model is not benefiting from knowledge distillation training when the teacher model is of the same capacity. We will revisit this problem in our discussion of out-of-domain results.
@@
We report the out-of-domain evaluation results in Table˜2. The baseline methods’ results are deferred to Appx.˜C.
We note similar observations as in the in-domain setting. Directly comparing contrastive learning to knowledge distillation, RankQwen-0.5B-KD achieves similar performance as RankQwen-0.5B-CL (53.1 average NDCG@10 v.s. 53.3); while RankQwen-1.5B-KD, RankQwen-3B-KD, RankRGemma-2B have better performance over their contrastive learning counterparts.
@@
Scaling model sizes.
Scaling has been proven effective in IR tasks (Neelakantan et al., 2022; Muennighoff et al., 2024b; Xu et al., 2025b; Zhu et al., 2023, inter alia). We also compare the model size scaling trend between CL versus KD for passage reranking in Fig.˜1. We make two observations:
  1. 1.
@@
Retriever Teacher Contrastive Learning Knowledge Distillation
Dataset 110M 7B 0.5B 3B 0.5B 3B
[…]
4.2 Scaling Training Data
Experiment setup.
In §˜4.1, we focus on a controlled experiment with various model sizes. Now we examine the efficacy of knowledge distillation when scaling the axis of training data. Notice that doing so leads to the decoupling of the retriever and reranker as we can no longer mine hard negatives from the retriever on MS MARCO passage trainset. We use a recently released trainset, RLHN Thakur et al. (2025), which consists of 649K $(q,d^{+},D^{-})$ training samples pruned from the larger BGE trainset Xiao et al. (2023). Specifically, RLHN comprises train samples from the following dataset: MS MARCO passage Bajaj et al. (2016), Arguana Wachsmuth et al. (2018), FEVER Thorne et al. (2018), FiQA Maia et al. (2018), SCIDOCS Cohan et al. (2020), HotpotQA Yang et al. (2018) and NQ Kwiatkowski et al. (2019). In this case, BEIR benchmark as a whole is no longer considered out-of-domain evaluation. We use a lightweight intfloat/e5-base-v2 retriever Wang et al. (2022b) pretrained on unlabeled text pairs, then finetuned with the combined MS MARCO passage and NQ training mixture.
Figure 2: Dataset scaling trend for CL versus KD. Left figure shows average results for 7 in-domain datasets from BEIR benchmark, while the right figure shows average results for 7 out-of-domain datasets.
We use the similar training strategy as §˜3 (the only difference is the ranker training data): we first train 7B-scale reranker with Qwen2.5-7B and contrastive learning, then train smaller rerankers with knowledge distillation, with the reranking scores on RLHN dataset labeled by the teacher, as well as the contrastive learning counterparts. We train 0.5B and 3B rerankers, as we notice Qwen2.5-1.5B fail to converge in the contrastive learning setting. We name the trained rerankers as RankQwen-{0.5B, 3B, 7B}-CL-RLHN and RankQwen-{0.5B, 3B}-KD-RLHN, respectively.
Results and analysis.
We report the results in Table˜3. Compared to Table˜2, we notice rerankers trained with RLHN mixture show significant performance boost, suggesting the effectiveness of in-domain training data.
Similar to Table˜2, the 7B model achieves the best overall performance, average 62.8 NDCG@10 on 14 BEIR benchmark datasets including MS MARCO Dev. We notice that knowledge distillation still outperforms contrastive learning when distilling from the strong 7B teacher model, though the margin is small (<1%). This observation suggests the robustness of knowledge distillation: it can achieve performance improvement compared to contrastive learning when the reranker training data is not coupled with the retriever.
In  Fig.˜2 we further analyze BEIR in-domain (7 datasets used in training) and out-of-domain (7 unseen datasets) performance with varying training data sizes. As the amount of training data increases, both CL and KD show improved in-domain performance, with CL eventually outperforming KD. In contrast, KD consistently outperforms CL on the OOD datasets, although the benefits of increasing training data are diminishing. We leave a more in-depth investigation of the in-domain versus OOD performance gap to future work.